OpenStudy (anonymous):

Find the slope of the line through the pair of points (9,-9) and (3,-7)

OpenStudy (anonymous):

Is -1/3 the answer?

OpenStudy (agent47):

M=(y-y1)/(x-x1)=(-9+7)/(9-3)=-2/6=-1/3. YOu are right

OpenStudy (agent47):

*You